Sketch the region enclosed by the given curves. Decide whether to integrate with respect to x or y. Then find the area of thecould you teach me how to do this two question?
Answer #1

vy=ufiy=2re g. 5x 24+2=5 73 solving 2n = S-u urnason utusa-5=0 ut58a-50-5=0 ( VI ) (Va+5)=0 = fuo x = || -5 Chin (veu) E (29—ཡང་ཞེར་ན་ཤི11 5 1² - Guto - 2 / yus (n = 3/2 + 9 y- =-567 32 y = 61-512 heet the line there onlsin beg=mn, which dirilder th____ye met - 6R-S 26- न pulting in 1072-3 (6-) to tox? 185 +15x²=36X3 2 53361 3633 12-5- - -9x3X4 725 - हाय 026-3.45 बाग--4)